OpenStudy (anonymous):

please how do i solve this 3log(2/t-2) - 2log(3/t-4) note that the t is raise to power -2 like wise -4 thanks

OpenStudy (anonymous):

To make sure I'm reading this right, it is 3*log(2/ (t raised to the power of -2)) - 2* log(3/(t to the power of -4)?

OpenStudy (anonymous):

\[3\times \ln (2\div t ^{-2}) - 2\times \ln (3\div t ^{-4})\]